OpenStudy (aravindg):

22. Let f(x) = ( x2 - 4x + 3, x<3 :: x-4, x≥ 3 ) and g(x) = ( x-3 , x < 4 x2 + 2 x + 2 , x≥ 4 ) determine f/g and find its domain

OpenStudy (anonymous):

f/g for x<3 is \[x ^{2}-4x+3=(x-1)(x-3)/x-3=x-1 for x<3\] and\[x-4/(x ^{2}+2x+2) for x \ge4\] i hope its correct

OpenStudy (aravindg):

othrs???